Write the equation of the line passing through the points(-1,6) and (2,0)
Answer by stanbon(75887)

Write the equation of the line passing through the points(-1,6) and (2,0)
----
slope = (6-0)/(-1-2) = 6/-3 = -2
-------------------------------------
Form: y = mx+b ; Solve for "b":
0 = -2*2 + b
b = 4
---------------
Equation:
y = -2x+4
===============
Cheers,
Stan H.